# Reinforcing Multimodal Understanding and Generation with Dual Self-rewards

### v2

@@ Abstract
Building upon large language models (LLMs), recent large multimodal models (LMMs) unify cross-model understanding and generation into a single framework. However, LMMs still struggle to achieve accurate image-text alignment, prone to generating text responses contradicting the visual input or failing to follow the text-to-image prompts. Current solutions require external supervision (e.g., human feedback or reward models) and only address unidirectional tasks—either understanding or generation. In this work, based on the observation that understanding and generation are inverse dual tasks, we introduce a self-supervised dual reward mechanism to reinforce the understanding and generation capabilities of LMMs. Specifically, we sample multiple outputs for a given input in one task domain, then reverse the input-output pairs to compute the dual likelihood of the model as self-rewards for optimization. Extensive experimental results on visual understanding and generation benchmarks demonstrate that our method can effectively enhance the performance of the model without any external supervision, especially achieving remarkable improvements in text-to-image tasks.
1 Introduction
@@
Recently, following the success of LLMs in text-centric tasks, many works have explored extending LLMs to multimodal tasks, leading to the emergence of LMMs. By optionally incorporating modality-specific encoders and decoders (or tokenizers and de-tokenizers), some LMMs are tailored for visual understanding tasks [1, 2], some for visual generation tasks [3], while others are designed to unify both understanding and generation tasks [4, 5, 6, 7].
@@
Despite the impressive performance of LMMs in both understanding and generation tasks, they still face challenges in aligning text and image modalities. For visual understanding tasks, LMMs often fail to accurately perceive visual content and return hallucinated or incorrect responses, as illustrated in LABEL:fig:example1. Ji et al. [8], Ouali et al. [9] conduct supervised fine-tuning (SFT) on paired data with preference labels. For text-to-image generation tasks, LMMs tend to generate images that are inconsistent with the text prompts, as shown in LABEL:fig:example2. Jiang et al. [10] propose a two-level chain-of-thought (COT) generation method and employ ensemble rule-based reward models to optimize the model. These efforts often concentrate on improving only one aspect (either understanding or generation) of LMM capabilities, and most of them depend heavily on external supervision signals or parallel text-image pairs data for training.
@@
We observe that understanding and generation are inherently dual tasks, where the output of one task can naturally serve as the input of the other. Moreover, existing unified LMMs have already undergone basic alignment and demonstrate a certain level of capability in both tasks, but the dual-task relationship is not fully utilized. Motivated by this, we seek to exploit this duality between understanding and generation to enable mutual self-supervision, allowing the optimization of one task to be guided by feedback derived from its dual counterpart. However, to derive self-supervised rewards for both tasks in a unified and efficient way is non-trivial. Existing methods tend to employ visual query-answer (VQA) tasks to assess the generated images in text-to-image (T2I) tasks [11, 12], which is not applicable to understanding tasks. DeGF [13] leverages an auxiliary image generated by Stable Diffusion [14] to improve the understanding performance, which relies on external feedback and is of high cost and latency.
@@
To achieve our objective and tackle the above challenges, we propose a dual self-reward mechanism to provide self supervision for optimizing the understanding and generation of LMMs. Specifically, for visual understanding, we sample multiple textual descriptions given an input image, then reverse the input-output pairs and use each description as the condition to compute the likelihood of generating the original image. This enables a quantitative assessment of how well each description aligns with the visual content. Analogously, for text-to-image generation, we compute the likelihood of generating the original text prompt conditioned on the sampled images to assess their semantic fidelity. By incorporating this dual self-reward mechanism, we can effectively optimize the understanding and generation capabilities in a unified manner. Besides, based on this dual self-reward mechanism, we explore different optimization strategies, including jointly optimizing both capabilities within a single model, as well as alternately optimizing the two capabilities in two adversarial models. We also investigate different optimization methods for this task, including preference optimization and reinforcement learning methods.
@@
Extensive experiments have been conducted on various visual understanding and generation benchmarks to validate the effectiveness of our method. Specifically, we evaluate the performance of our method on the T2I-CompBench [15] and GenEval [16] datasets, which are widely used for assessing the performance of LMMs in text-to-image generation tasks. Our tuned Janus-Pro-7B [7] model gains remarkable improvements over the baseline model, achieving an 11.68% increase on average on T2I-CompBench and a 5% increase on GenEval. Additionally, the model also demonstrates improved performance on visual understanding tasks, for example, achieving a 7.5% increase in the overall score on the LLaVA-Bench  [^1] https://huggingface.co/datasets/liuhaotian/llava-bench-in-the-wild.
In summary, our contributions are as follows:
@@
  • To the best of our knowledge, we present the first unified self-supervised framework that jointly enhances both the understanding and generation capabilities of large multimodal models (LMMs).
@@
  • We propose a dual self-reward mechanism that leverages the inherent duality between understanding and generation tasks to provide self-supervised optimization signals. And we explore different optimization strategies and optimization methods based on the dual self-rewards for unified learning of LMMs.
@@
2 Related Work
2.1 Multimodal Understanding and Generation
@@
Advancement in large language models (LLMs) has paved the way for the development of large multimodal models (LMMs) for vision-language understanding. These works focus on how to align visual signals with the language semantic space of the pre-trained LLMs. By connecting vision encoders to pre-trained LLMs, LMMs can effectively understand both visual and textual inputs [17, 18]. LLaVA [2] and MiniGPT-4 [19] were among the first to attempt to train adapters to shift the output of the vision encoder into the embedding space of the language model. Qwen-VL [1], InternVL [20] and Deepseek-VL [21, 22] further enhance the visual-linguistic performance on more complex scenarios by incorporating more advanced LLM backbones, additional training data of higher quality and more effective training strategy. Although these models have achieved impressive performance in visual understanding and reasoning, mere perception and comprehension fall short of our broader expectations for multimodal intelligence.
@@
Recent advances in visual generation can be generally categorized into two main approaches: diffusion models and autoregressive models. The LMMs for visual generation fall in the latter category, which we mainly focus on in this paper. These models typically adopt an autoregressive approach, which is the focus of this paper. These models generate images by producing a sequence of discrete visual tokens, which are then decoded into images using vector-quantized (VQ) model [23, 24] based visual tokenizers. Representative works in this line of research include LlamaGen and DeLVM, which demonstrate the potential of autoregressive modeling in high-fidelity image synthesis within a multimodal framework.
@@
Unified LMMs aim to integrate both understanding and generation capabilities into a single framework, enabling them to perform a wide range of multimodal tasks. Unified LMMs aim to seamlessly integrate understanding and generation capabilities within a single framework, enabling them to handle a broad spectrum of multimodal tasks. Models like Emu [25], Emu2 [26], X-ViLa [27], and Next-GPT [28] adopt a unified autoregressive architecture that treats both visual embeddings and text tokens as elements to predict. AnyGPT [29], Show-o [5], ViLA-U [30], Chameleon [31], Emu3[4], and Janus series [6, 7], leverage VQ tokenizers to discretize images into tokens, thereby allowing both vision and language to be modeled uniformly through next-token prediction.
2.2 Optimization for Large Multimodal Models
Although the LMMs demonstrated promising performance in various multimodal tasks, they still face challenges in accurately aligning textual and visual modalities. Existing works explore how to conduct fine-grained optimization on LMMs to further improve them. Align-anything [8] propose a framework to align all-modality (any-to-any) models with human intentions using unified language feedback and 200k human preference annotations. SILMM [11] proposes a self-VQA manner and DPO [32] to improve the compositional text-to-image generation performance, while T2I-R1 utilizes ensemble rule-based reward models to reinforce two levels of chain-of-thought (COT) text-to-image generation with GRPO [33]. These methods focus on improving the performance of LMMs in either understanding or generation tasks, and most of them rely on external supervision signals, such as human feedback or reward models, which may not be readily available in all scenarios. Based on the observation that understanding and generation are inherently dual tasks and they can assess the output of each other, we explore a self-supervised dual reward mechanism to reinforce the bi-directional tasks in a unified manner.
Figure 2: Overview of our dual self-reward mechanism. The model samples multiple outputs for a given input for one task and reverses the input-output pairs to compute the dual likelihoods of the model outputs conditioned on the original input as self-rewards for optimization.
3 Methodology
In this section, we first introduce the duality between understanding and generation tasks in unified LMMs. Then we present our dual self-reward mechanism, which leverages the inherent duality between understanding and generation tasks to provide self-supervised optimization signals. Finally, we describe the optimization method used for model optimization based on the self-reward signals. The overview of our dual self-reward mechanism is shown in Figure 2
3.1 Duality between Understanding and Generation
@@
$\pi_{\theta}(Y_{V}\mid X_{T})=\prod_{i=1}^{\|Y_{V}\|}\pi_{\theta}(y_{v}^{i}\mid X_{T},y_{v}^{<i}),$ (1)
@@
Reversely, for visual understanding, the model takes an image $X_{V}$ as input and generates a text response $Y_{T}=\{y_{t}^{i}\}_{i=1}^{\|Y_{T}\|}$ that matches the content of the image. This process can be formulated as:
$\pi_{\theta}(Y_{T}\mid X_{V})=\prod_{i=1}^{\|Y_{T}\|}\pi_{\theta}(y_{t}^{i}\mid X_{V},y_{t}^{<i}),$ (2)
@@
3.2 Dual Self-reward Mechanism
@@
$R_{U}(Y_{T}\mid X_{V})$ $=\frac{1}{\|X_{V}\|}\log\pi_{\theta}(X_{V}\mid Y_{T})=\frac{1}{\|X_{V}\|}\sum_{1}^{\|X_{V}\|}\log\pi_{\theta}(x_{v}^{i}\mid Y_{T},x_{v}^{<i}),$ (3)
$R_{G}(Y_{V}\mid X_{T})$ $=\frac{1}{\|X_{T}\|}\log\pi_{\theta}(X_{T}\mid Y_{V})=\frac{1}{\|X_{T}\|}\sum_{1}^{\|X_{T}\|}\log\pi_{\theta}(x_{t}^{i}\mid Y_{V},x_{t}^{<i}),$
@@
Our dual self-reward mechanism naturally possesses the following advantages: 1) efficiency, compared to self-VQA and generating additional reference content, dual self-rewards can be computed in a single forward pass of the model; 2) self-supervised, the dual self-rewards are derived from the model’s own outputs, eliminating the need for external supervision or additional data; 3) unified, the dual self-rewards can be applied to both understanding and generation tasks, allowing for a more comprehensive optimization process; 4) unbiased, the rewards are likelihoods of the same original input for multiple sampled outputs, which can reduce the length bias [34] existing in the reward model based methods [8].
3.3 Optimization with Self-rewards
With the dual self-rewards, we can optimize the model using different optimization methods, including preference optimization and reinforcement learning methods, as well as different optimization strategies.
We adopt the online SimPO [35] method, which is a simple preference optimization method that can be easily applied to LMMs. We also employ the GRPO [33] method, which is a prevalent reinforcement learning method to optimize the reasoning capabilities of LLMs.
@@ 3.3 Optimization with Self-rewards
$\left[\log\sigma\left(\frac{\beta}{\|Y^{+}\|}\log\pi_{\theta}(Y^{+}\mid X)-\frac{\beta}{\|Y^{-}\|}\log\pi_{\theta}(Y^{-}\mid X)-\gamma\right)\right],$ (4)
@@
We also employ the GRPO [33] method, which estimates the advantage in a group-relative manner without relying on a value function, making it suitable for optimizing LMMs with self-reward signals. Following DAPO [36], we adopt a higher clipping threshold and a token-level policy gradient loss in the GRPO objective.
@@
$\hat{A}_{i,t}=\frac{R^{(i)}-\text{mean}(\{R^{(j)}\}_{j=1}^{G})}{\text{std}(\{R^{(j)}\}_{j=1}^{G})}.$ (5)
@@
$r_{i,t}(\theta)=\frac{\pi_{\theta}(y^{(i)}_{t}\mid X,y^{(i)}_{<t})}{\pi_{\theta_{\text{ref}}}(y^{(i)}_{t}\mid X,y^{(i)}_{<t})}.$ (6)
@@
$\left[\frac{1}{\sum_{i=1}^{G}\|Y^{(i)}\|}\sum_{i=1}^{G}\sum_{t=1}^{\|Y^{(i)}\|}\left(\min\left(\begin{aligned} &r_{i,t}(\theta)\hat{A}_{i,t},\\ &\text{clip}(r_{i,t}(\theta),1-\epsilon_{\text{low}},1+\epsilon_{\text{high}})\hat{A}_{i,t}\end{aligned}\right)-\beta D_{\text{KL}}(\pi_{\theta}\parallel\pi_{\text{ref}})\right)\right],$ (7)
@@
4 Experiments
4.1 Experimental Setup

### v3

# SUDER: Self-Improving Unified Large Multimodal Models for Understanding and Generation with Dual Self-Rewards

@@ Abstract.
Building upon large language models (LLMs), recent large multimodal models (LMMs) unify cross-model understanding and generation into a single framework. However, LMMs still struggle to achieve accurate vision-language alignment, prone to generating text responses contradicting the visual input or failing to follow the text-to-image prompts. Current solutions require external supervision (e.g., human feedback or reward models) and only address unidirectional tasks—either understanding or generation. In this work, based on the observation that understanding and generation are naturally inverse dual tasks, we propose SUDER (Self-improving Unified LMMs with Dual sElf-Rewards), a framework reinforcing the understanding and generation capabilities of LMMs with a self-supervised dual reward mechanism. SUDER leverages the inherent duality between understanding and generation tasks to provide self-supervised optimization signals for each other. Specifically, we sample multiple outputs for a given input in one task domain, then reverse the input-output pairs to compute the dual likelihood within the model as self-rewards for optimization. Extensive experimental results on visual understanding and generation benchmarks demonstrate that our method can effectively enhance the performance of the model without any external supervision, especially achieving remarkable improvements in text-to-image tasks.
Large Multimodal Model, Text-to-image, Vision Language Model, Self-Supervised Learning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Artificial intelligence
1. Introduction
@@
Recently, following the success of LLMs in text-centric tasks, many works have explored extending LLMs to multimodal tasks, leading to the emergence of LMMs. By optionally incorporating modality-specific encoders and decoders (or tokenizers and de-tokenizers), some LMMs are tailored for visual understanding tasks (Bai et al., 2023; Liu et al., 2023), some for visual generation tasks (Sun et al., 2024b), while others are designed to unify both understanding and generation tasks (Wang et al., 2024b; Xie et al., 2024; Wu et al., 2024b; Chen et al., 2025).
@@
Despite the impressive performance of LMMs in both understanding and generation tasks, they still face challenges in aligning text and image modalities. For visual understanding tasks, LMMs often fail to accurately perceive visual content and return hallucinated or incorrect responses, as illustrated in Figure 1(a). Ji et al. (2024); Ouali et al. (2024) conduct supervised fine-tuning (SFT) on paired data with preference labels to mitigate this problem. For text-to-image generation tasks, LMMs tend to generate images that are inconsistent with the text prompts, as shown in Figure 1(b). Jiang et al. (2025) propose a two-level chain-of-thought (COT) generation method and employ ensemble rule-based reward models to optimize the model. Wang et al. (2025) utilize CLIP (Radford et al., 2021) as the reward model to supervise the reinforcement learning of generation. These efforts often concentrate on improving only one aspect (either understanding or generation) of LMM capabilities, and most of them depend heavily on external supervision signals or parallel text-image pairs data for training.
@@
(a) Misalignment in visual understanding.
(b) Misalignment in text-to-image generation.
Figure 1. Failures in visual understanding and image generation of Janus-Pro-7B. In (a), the text in $\underline{\text{red}}$ indicates error content. In (b), the generated images do not follow the content in orange.
@@
We observe that understanding and generation are inherently dual tasks, where the output of one task can naturally serve as the input of the other. Moreover, existing unified LMMs have already undergone basic alignment and demonstrate a certain level of capability in both tasks, but the dual-task relationship is not fully utilized. Motivated by this, we seek to exploit this duality between understanding and generation to enable mutual self-supervision, allowing the optimization of one task to be guided by feedback derived from its dual counterpart. By this, we can achieve self-improvement of the unified LMMs in both understanding and generation tasks without the need for external human or AI supervision. However, to derive self-supervised rewards for both tasks in a unified and efficient way is non-trivial. Existing methods tend to employ visual query-answer (VQA) tasks to assess the generated images in text-to-image (T2I) tasks (Qu et al., 2025; Wang et al., 2024a), which are not applicable to understanding tasks. DeGF (Zhang et al., 2025) leverages an auxiliary image generated by Stable Diffusion (Rombach et al., 2022) to improve the understanding performance, which relies on feedback from external models and is of high cost and latency.
@@
To achieve our objective and tackle the above challenges, we propose a dual self-reward mechanism to provide self supervision for optimizing the understanding and generation of LMMs. Specifically, for visual understanding, we sample multiple textual descriptions given an input image, then reverse the input-output pairs and use each description as the condition to compute the likelihood of generating the original image. This enables a quantitative assessment of how well each description aligns with the visual content. Analogously, for text-to-image generation, we compute the likelihood of generating the original text prompt conditioned on the sampled images to assess their semantic fidelity. By incorporating this dual self-reward mechanism, we introduce a framework self-improving unified LMMs SUDER (Self-improving Unified LMMs with Dual sElf-Rewards), which can effectively optimize the understanding and generation capabilities in a unified manner. Besides, we explore different optimization strategies, including jointly optimizing both capabilities within a single model, as well as alternately optimizing the two capabilities in two adversarial models. We also investigate different optimization algorithms for this task, including preference optimization and reinforcement learning algorithms.
@@
Extensive experiments have been conducted on various visual understanding and generation benchmarks to validate the effectiveness of our method. Specifically, we evaluate the performance of our method on the T2I-CompBench (Huang et al., 2023) and GenEval (Ghosh et al., 2023) datasets, which are widely used for assessing the performance of LMMs in text-to-image generation tasks. For example, our tuned Janus-Pro-7B (Chen et al., 2025) model gains remarkable improvements over the baseline model, achieving an 11.68% increase on average on T2I-CompBench and a 5% increase on GenEval. Additionally, the model also demonstrates improved performance on visual understanding tasks, for example, achieving a 7.5% increase in the overall score on the LLaVA-Bench  [^1] https://huggingface.co/datasets/liuhaotian/llava-bench-in-the-wild.
Overall, our contributions can be summarized as follows:
@@
  • We present SUDER, a unified self-supervised framework that jointly enhances both the understanding and generation capabilities of large multimodal models (LMMs).
@@
  • We propose a dual self-reward mechanism that leverages the inherent duality between understanding and generation tasks to provide self-supervised optimization signals. And we explore different training strategies and optimization algorithms based on the dual self-rewards for unified learning of LMMs.
@@
2. Related Work
2.1. Multimodal Understanding and Generation
@@
Advancement in large language models (LLMs) has paved the way for the development of large multimodal models (LMMs) for vision-language understanding. These works focus on how to align visual signals with the language semantic space of the pre-trained LLMs. By connecting vision encoders to pre-trained LLMs, LMMs can effectively understand both visual and textual inputs (OpenAI, 2023; Anthropic, 2024). LLaVA (Liu et al., 2023) and MiniGPT-4 (Zhu et al., 2024) were among the first to attempt to train adapters to shift the output of the vision encoder into the embedding space of the language model. Qwen-VL (Bai et al., 2023), InternVL (Chen et al., 2024b) and Deepseek-VL (Lu et al., 2024; Wu et al., 2024a) further enhance the visual-linguistic performance on more complex scenarios by incorporating more advanced LLM backbones, additional training data of higher quality and more effective training strategy. Although these models have achieved impressive performance in visual understanding and reasoning, mere perception and comprehension fall short of our broader expectations for multimodal intelligence.
@@
Recent advances in visual generation can be generally categorized into two main approaches: diffusion models and autoregressive models. The LMMs for visual generation fall in the latter category, which we mainly focus on in this paper. These models typically adopt an autoregressive approach, which is the focus of this paper. These models generate images by producing a sequence of discrete visual tokens, which are then decoded into images using vector-quantized (VQ) model (van den Oord et al., 2017; Esser et al., 2021) based visual tokenizers. Representative works in this line of research include LlamaGen and DeLVM, which demonstrate the potential of autoregressive modeling in high-fidelity image synthesis within a multimodal framework.
@@
Unified LMMs aim to integrate both understanding and generation capabilities into a single framework, enabling them to perform a wide range of multimodal tasks. Unified LMMs aim to seamlessly integrate understanding and generation capabilities within a single framework, enabling them to handle a broad spectrum of multimodal tasks. Models like Emu (Sun et al., 2024c), Emu2 (Sun et al., 2024a), X-ViLa (Ye et al., 2024), and Next-GPT (Wu et al., 2024c) adopt a unified autoregressive architecture that treats both visual embeddings and text tokens as elements to predict. AnyGPT (Zhan et al., 2024), Show-o (Xie et al., 2024), ViLA-U (Wu et al., 2024d), Chameleon (Lu et al., 2023), Emu3(Wang et al., 2024b), and Janus series (Wu et al., 2024b; Chen et al., 2025), leverage VQ tokenizers to discretize images into tokens, thereby allowing both vision and language to be modeled uniformly through next-token prediction.
2.2. Optimization for Large Multimodal Models
Although the LMMs demonstrated promising performance in various multimodal tasks, they still face challenges in accurately aligning textual and visual modalities. Existing works explore how to conduct fine-grained optimization on LMMs to further improve them. Align-anything (Ji et al., 2024) propose a framework to align all-modality (any-to-any) models with human intentions using unified language feedback and 200k human preference annotations. SILMM (Qu et al., 2025) proposes a self-VQA manner and DPO (Rafailov et al., 2023) to improve the compositional text-to-image generation performance, while T2I-R1 utilizes ensemble rule-based reward models to reinforce two levels of chain-of-thought (COT) text-to-image generation with GRPO (Shao et al., 2024). These methods focus on improving the performance of LMMs in either understanding or generation tasks, and most of them rely on external supervision signals, such as human feedback or reward models, which may not be readily available in all scenarios. Based on the observation that understanding and generation are inherently dual tasks and they can assess the output of each other, we explore a self-supervised dual reward mechanism to reinforce the bi-directional tasks in a unified manner.
Figure 2. Overview of SUDER. The model samples multiple outputs for a given input for one task and reverses the input-output pairs to compute the dual likelihoods of the model outputs conditioned on the original input as self-rewards, which are utilized in optimization algorithms.
3. Methodology
In this section, we first introduce the duality between understanding and generation tasks in unified LMMs. Then we present our dual self-reward mechanism, which leverages the inherent duality between understanding and generation tasks to provide self-supervised optimization signals. Finally, we describe the framework SUDER, which incorporates the dual self-reward mechanism and explores different optimization strategies and algorithms for unified learning of LMMs. The overview of SUDER is shown in Figure 2.
3.1. Duality between Understanding and Generation
@@
(1) $\pi_{\theta}(Y_{V}\mid X_{T})=\prod_{i=1}^{\|Y_{V}\|}\pi_{\theta}(y_{v}^{i}\mid X_{T},y_{v}^{<i}),$
@@
Reversely, for visual understanding, we take the captioning task as the proxy. The model takes an image $X_{V}$ as input and generates a text response $Y_{T}=\{y_{t}^{i}\}_{i=1}^{\|Y_{T}\|}$ that matches the content of the image. This process can be formulated as:
(2) $\pi_{\theta}(Y_{T}\mid X_{V})=\prod_{i=1}^{\|Y_{T}\|}\pi_{\theta}(y_{t}^{i}\mid X_{V},y_{t}^{<i}),$
@@
3.2. Dual Self-reward Mechanism
@@
(3) $R_{U}(Y_{T}\mid X_{V})$ $=\frac{1}{\|X_{V}\|}\log\pi_{\theta}(X_{V}\mid Y_{T})$
$=\frac{1}{\|X_{V}\|}\sum_{1}^{\|X_{V}\|}\log\pi_{\theta}(x_{v}^{i}\mid Y_{T},x_{v}^{<i}),$
$R_{G}(Y_{V}\mid X_{T})$ $=\frac{1}{\|X_{T}\|}\log\pi_{\theta}(X_{T}\mid Y_{V})$
$=\frac{1}{\|X_{T}\|}\sum_{1}^{\|X_{T}\|}\log\pi_{\theta}(x_{t}^{i}\mid Y_{V},x_{t}^{<i}),$
@@
Our dual self-reward mechanism naturally possesses the following advantages: 1) efficiency, compared to self-VQA and generating additional reference content, dual self-rewards can be computed in a single forward pass of the model; 2) self-supervised, the dual self-rewards are derived from the model’s own outputs, eliminating the need for external supervision or additional data; 3) unified, the dual self-rewards can be applied to both understanding and generation tasks, allowing for a more comprehensive optimization process; 4) unbiased, the rewards are likelihoods of the same original input for multiple sampled outputs, which can reduce the length bias (Shen et al., 2023) existing in the reward model based methods (Ji et al., 2024).
3.3. SUDER: Self-improving ULMMs with DSR
SUDER is a framework that incorporates the dual self-reward mechanism to the optimization process. With the dual self-rewards, we can optimize the model using different optimization algorithms, including preference optimization and reinforcement learning algorithms, as well as different training strategies.
We adopt the online SimPO (Meng et al., 2024) method, which is a simple preference optimization method that can be easily applied to LMMs. We also employ the GRPO (Shao et al., 2024) method, which is a prevalent reinforcement learning method to optimize the reasoning capabilities of LLMs.
@@ 3.3. SUDER: Self-improving ULMMs with DSR
[…]
(4) $\quad\left.\left.-\frac{\beta}{\|Y^{-}\|}\log\pi_{\theta}(Y^{-}\mid X)-\gamma\right)\right].$
@@
We also employ the GRPO (Shao et al., 2024) method, which estimates the advantage in a group-relative manner without relying on a value function, making it suitable for optimizing LMMs with self-reward signals. Following DAPO (Yu et al., 2025), we adopt a higher clipping threshold and a token-level policy gradient loss in the GRPO objective.
@@
(5) $\hat{A}_{i,t}=\frac{R^{(i)}-\text{mean}(\{R^{(j)}\}_{j=1}^{G})}{\text{std}(\{R^{(j)}\}_{j=1}^{G})}.$
@@
(6) $r_{i,t}(\theta)=\frac{\pi_{\theta}(y^{(i)}_{t}\mid X,y^{(i)}_{<t})}{\pi_{\theta_{\text{ref}}}(y^{(i)}_{t}\mid X,y^{(i)}_{<t})}$.
@@
$\left[\frac{1}{\sum_{i=1}^{G}\|Y^{(i)}\|}\sum_{i=1}^{G}\sum_{t=1}^{\|Y^{(i)}\|}\left(\texttt{minclip}(r_{i,t}(\theta)\hat{A}_{i,t})-\beta D_{\text{KL}}(\pi_{\theta}\parallel\pi_{\text{ref}})\right)\right],$
(7) $\texttt{minclip}(r_{i,t}(\theta)\hat{A}_{i,t})=\min\left(\begin{aligned} &r_{i,t}(\theta)\hat{A}_{i,t},\\ &\text{clip}(r_{i,t}(\theta),1-\epsilon_{\text{low}},1+\epsilon_{\text{high}})\hat{A}_{i,t}\end{aligned}\right),$
@@
4. Experiments
4.1. Experimental Setup